# 3DTopia-XL: Scaling High-quality 3D Asset Generation via Primitive Diffusion

### v1

@@ 4.3 Image-to-3D Generation
In this section, we compare our single-view conditioned generative model with state-of-the-art methods suitable for image-to-3D synthesis.
Comparison Methods.
@@
We run evaluations against two types of methods: 1) sparse-view reconstruction models, and 2) image-conditioned diffusion models. The reconstruction-based models, like LGM (Tang et al., 2024), InstantMesh (Xu et al., 2024a), Real3D (Jiang et al., 2024), CRM (Wang et al., 2024), are deterministic methods that learn to reconstruct 3D objects given four or six input views. They enable single-view to 3D synthesis by leveraging pretrained diffusion models (Shi et al., 2023; Li et al., 2024b) to generate multiple views from the input single image. However, as methods for reconstruction heavily rely on the input multi-view images, those methods suffer from multi-view inconsistency caused by the frontend 2D diffusion models. The feed-forward diffusion models, like CraftsMan (Li et al., 2024c), Shap-E (Jun & Nichol, 2023), LN3Diff (Lan et al., 2024), are probabilistic methods that learn to generate 3D objects given input image conditions. All methods above only model the shape and color without considering roughness and metallic while our method is suitable to produce those assets.
Results.
@@
Figure 6 demonstrates qualitative results. To fairly compare the capability of generating 3D assets ready for rendering, we take the exported textured mesh from each method into Blender (Community, 2018) and render it with the target environment map. For methods that cannot produce PBR materials, we assign the default diffuse material. Existing reconstruction-based models fail to produce good results which may suffer from multiview inconsistency and incapability to support spatially varied materials. Moreover, these reconstruction models are built upon triplane representation which is not parameter-efficient. This downside limits the spatial resolution of the underlying 3D representation, leading to the bumpy surface indicated by the rendered normal. On the other hand, existing 3D diffusion models fail to generate objects that are visually aligned with the input condition. While CraftsMan is the only method that has comparable surface quality as ours, they are only capable of generating 3D shapes without textures and materials. In contrast, 3DTopia-XL achieves the best visual and geometry quality among all methods. Thanks to our capability to generate spatially varied PBR assets (metallic/roughness), our generated mesh can also produce vivid reflectance with specular highlights even under harsh environmental illuminations.
User Study.
We conduct an extensive user study to quantitatively evaluate image-to-3D performance. We opt for an output evaluation (Bylinskii et al., 2022) for user study, where each volunteer is shown with a pair of results comparing a random method against ours, and asked to choose the better one in four aspects: 1) Overall Quality, 2) Image Alignment, 3) Surface Smoothness, and 4) Physical Correctness. A total number of 48 paired samples are provided to 27 volunteers for the flip test. We summarize the average preference percentage across all four dimensions in Figure 7. 3DTopia-XL is the best one among all methods. Although the image alignment of our method is only a slight improvement against reconstruction-based methods like CRM, the superior quality of geometry and the ability to model physically based materials are the keys to producing the best overall quality in the final rendering.
Figure 7: User Study. We quantitatively evaluate comparison methods by conducting preference tests against our method on four dimensions. The results show that 3DTopia-XL has the highest preference rate compared with each of other methods.
4.4 Text-to-3D Generation
We also demonstrate our capability for native text-to-3D generation as shown in Figure 4. Note that, as a 3D native diffusion model, our text-driven generation is done by directly conditioning the model on textual inputs, without relying on complicated text-to-multiview followed by reconstruction models.
Moreover, we also conduct quantitative evaluations against native text-to-3D generative models. Given a set of unseen text prompts, we take the CLIP Score as the evaluation metric which is the cosine similarity between the text embedding and image embedding in the joint text-image space of the CLIP model (Radford et al., 2021). We take the front-view rendering from each method to compute the image embedding. We mainly compare two methods with open-source implementations: Shap-E (Jun & Nichol, 2023) and 3DTopia (Hong et al., 2024). Shap-E directly generates implicit functions of 3D objects conditioned on texts. 3DTopia adopts a hybrid 2D and 3D diffusion prior by using feedforward triplane diffusion followed by optimization-based refinement. As shown in Table 5, our method achieves better alignment between input text and rendering of generated asset.
4.5 Further Analysis
4.5.1 Number and Resolution of Primitives
As a structured 3D representation and a serialized representation, the number of primitives $N$ and the resolution of each primitive $a$ are two critical factors for the efficiency-quality tradeoff in PrimX. More and larger primitives often lead to better approximation quality. However, it results in a longer sequence length and deeper feature dimensions, causing inefficient long-context attention computation and training difficulty of the diffusion model. Therefore, we explore the impact of the number and resolution of primitives on different parameter budgets. We evaluate the PSNR of SDF, albedo, and material values given 500k points sampled near the surface. As shown in Table 2, given a fixed parameter count, a longer sequence of primitives appears to have a better approximation of SDF, texture, and material. Moreover, increasing the resolution of each primitive can reduce the approximation error. However, its benefit is marginal as the number of primitives is enough. The visualization in Figure 8 also confirms this observation. The alternative with $(N=64,a=32)$ produces poor geometry even with more parameter count since larger local primitives have higher chances to waste parameters in empty space. Furthermore, a longer sequence will increase the GFlops of DiT which also leads to better generation quality (Table 3). Therefore, we tend to use a longer sequence of primitives with a relatively small local resolution. Note that, the compression rate of our VAE (Sec. 3.2) also influences the hyperparameter of PrimX, which we will investigate in the next section.
@@ 4.5.1 Number and Resolution of Primitives
Figure 8: Ablation studies of the number and resolution of primitives. Our final setting $(N=2048,a=8)$ has the optimal approximation quality of ground truth, especially for fine-grained details like thin rotor blades. We visualize the corresponding PrimX at the bottom.
Table 2: Quantitative analysis of the number ($N$) and resolution ($a$) of primitives.
@@
4.5.2 Patch Compression Rate
The compression rate of our primitive patch-based VAE is also an important design choice. Overall, as a patch-based compression, we aim to do spatial compression of each primitive to save computation instead of semantic compression (Rombach et al., 2022). Empirically, a higher compression rate leads to a more efficient latent diffusion model, indicating larger batch sizes or larger models when scaling up. On the contrary, extreme compression often accompanies loss of information. Therefore, we analyze different compression rates given two different sequence lengths $N=256$ and $N=2048$ with the same parameter count of PrimX. For the evaluation metric, we compute the PSNR between the VAE’s output and input on 1k random samples from the dataset to measure its reconstruction quality. Table 5 shows the results where the final choice of $N=2048$ with compression rate $f=48$ achieves the optimal VAE reconstruction. The setting with $N=256,f=48$ has the same compression rate but lower reconstruction quality and a latent space with higher resolution, which we find difficulty in the convergence of the latent primitive diffusion model $g_{\Phi}$.
Figure 9: Scaling up 3DTopia-XL improves FID.
Setting Rendering-FID $\downarrow$
$N=256$ 76.31
$N=2048$ 16.16
Table 4: Analysis of different compression rates for VAE. $f$ stands for the compression rate between input and latent.
@@ 4.5.2 Patch Compression Rate
4.5.3 Scaling
@@
We further investigate the scaling law of 3DTopia-XL against model sizes and iterations. For metrics, we use Fréchet Inception Distance (FID) computed over 5k random samples without CFG guidance. Specifically, we consider Latent-FID which is computed in the latent space of our VAE and Rendering-FID which is computed on the DINO (Oquab et al., 2023) embeddings extracted from images rendered with Eq. 6. Figure 9 shows how Latent-FID and Rendering-FID change as the model size increases. We observe consistent improvements as the model becomes deeper and wider. Table 3 also demonstrates that longer sequence (smaller patches) leads to better performance, which may come from the findings in the vanilla DiT that increasing GFlops leads to better performance.
4.5.4 Sampling Diversity

### v2

@@ 4.2 Image-to-3D Generation
Comparison Methods. We run evaluations against two types of methods: 1) sparse-view reconstruction models, and 2) image-conditioned diffusion models. The reconstruction-based models, like LGM [56], InstantMesh [68], Real3D [20], CRM [62], are deterministic methods that learn to reconstruct 3D objects given four or six input views. They enable single-view to 3D synthesis by leveraging pretrained diffusion models [52, 29] to generate multiple views from the input single image. However, those methods may suffer from multi-view inconsistency caused by the frontend 2D diffusion models. The feed-forward diffusion models, like CraftsMan [30], Shap-E [21], LN3Diff [26], are probabilistic methods that learn to generate 3D objects given input image conditions. All methods above only model the shape and color without considering roughness and metallic while our method is suitable to produce those assets with sampling diversity. To fairly compare the quality in real-world use cases, we place the output mesh of each method into Blender [10] and render it with an environment map. For methods that cannot produce PBR materials, we assign the default diffuse material.
@@
Results. Figure 5 demonstrates qualitative results. Existing reconstruction-based models fail to produce good results which may suffer from multiview inconsistency and incapability to support spatially varied materials. Moreover, most of reconstruction models are built upon triplane representation which is not parameter-efficient. This downside limits the spatial resolution of the underlying 3D representation, leading to the bumpy surface indicated by the rendered normal. On the other hand, existing 3D diffusion models fail to generate objects that are visually aligned with the input condition. While CraftsMan is the only method with a comparable surface quality as ours, they are only capable of generating 3D shapes without textures and materials. In contrast, 3DTopia-XL achieves the best visual and geometry quality among all methods. Thanks to our capability to generate spatially varied PBR assets, our generated mesh can produce vivid reflectance with specular highlights even under harsh environmental illuminations.
@@
Figure 6: Ablation studies of the number and resolution of primitives. Our final setting $(N=2048,a=8)$ has the optimal approximation quality of ground truth, especially for fine-grained details like thin rotor blades.
Table 2: Analysis of number ($N$) and resolution ($a$) of PrimX.
@@
4.3 Text-to-3D Generation
We conduct quantitative evaluations against native text-to-3D generative models, which achieves text-to-3D generation without text-to-multiview diffusion models. Given a set of unseen text prompts, we take the CLIP Score [48] as the evaluation metric following previous work [21, 18]. We mainly compare two methods with open-source implementations: Shap-E [21] and 3DTopia [18]. As shown in Table 4, our method achieves better alignment between input text and rendering of the generated asset. We defer the qualitative results in the supplementary.
4.4 Further Analysis
Number and Resolution of Primitives. As a structured and serialized 3D representation, the number of primitives and the resolution of each primitive are two critical factors for the efficiency-quality tradeoff in PrimX. More and larger primitives often lead to better approximation quality. However, it results in a longer set length and deeper feature dimensions, causing inefficient long-context attention computation and training difficulty of the diffusion model. Therefore, we explore the impact of the number and resolution of primitives on different parameter budgets. We evaluate the PSNR of SDF, albedo, and material values given 500k points sampled near the surface. As shown in Table 2 and Figure 6, given a fixed parameter count, a larger set with compact primitives lead to better approximations. Furthermore, a longer sequence will increase the GFlops of DiT which also leads to better generation quality shown in the supplementary. Therefore, we use a large set of primitives with a relatively small local resolution.
Patch Compression Rate. The spatial compression rate of our primitive patch-based VAE (Sec. 3.2) is also an important design choice. Empirically, a higher compression rate leads to a more efficient latent diffusion model but risks information loss. Therefore, we analyze different compression rates given two different set lengths $N=256$ and $N=2048$ with the same parameter count of PrimX. We measure the PSNR between the VAE’s output and input on 1k random samples. Table 4 shows the results where the final choice of $N=2048$ with compression rate $f=48$ achieves the optimal VAE reconstruction. The setting with $N=256,f=48$ has the same compression rate but lower reconstruction quality and a latent space with higher resolution, which we find difficulty in the convergence of the latent primitive diffusion model $g_{\Phi}$.
Inpainting and Interpolation. We show 3D generative applications like 3D inpainting and interpolation in Figure 7, showing the unique property of 3D native diffusion model compared to reconstruction methods.
Besides the ablation studies above, we also analyze 1) the model scaling, 2) the sampling diversity, 3) PrimX initialization, 4) user study, 5) more comparisons and 6) more visual results, which are deferred to the supplementary due to the space limit.
Table 3: Analysis of different compression rates for VAE. The scalar $f$ stands for the compression rate between the input to the VAE and the latent code (i.e. the ratio between the 2nd and 3rd column of the table).
@@ 4.4 Further Analysis
Figure 7: 3D Generative Applications. a) 3D Inpainting for image-to-3D and b) Interpolation between text-to-3D samples.
5 Conclusion
We present 3DTopia-XL, a native 3D diffusion model for PBR asset generation given textual or visual inputs. Central to our approach is PrimX, an innovative primitive-based 3D representation that is parameter-efficient, tensorial, and renderable. It encodes shape, albedo, and material into a compact $N\times D$ tensor, enabling the modeling of high-resolution geometry with PBR assets. On top of PrimX, we propose Latent Primitive Diffusion for scalable 3D generative models, together with practical techniques to export PBR assets ready for graphics pipelines.
Acknowledgement. This study is supported by the National Key R&D Program of China No.2022ZD0160102. This research/project is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG2-PhD-2021-08-019), the Ministry of Education, Singapore, under its MOE AcRF Tier 2 (MOET2EP20221- 0012, MOE-T2EP20223-0002), NTU NAP, and under the RIE2020 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) Funding Initiative, as well as cash and in-kind contribution from the industry partner(s).
Appendix A Appendix
This supplementary material is organized as follows:
  • Sec. A.1 provides further discussions, including the main difference between PrimX and existing 3D representations (Sec. A.1.1) and limitations (Sec. A.1.2).
  • Sec. A.2 introduces further experiments and evaluations, including quantitative results on Objaverse [11] and GSO [12] datasets (Sec. A.2.1), user study (Sec. A.2.3), model scaling (Sec. A.2.4), sampling diversity (Sec. A.2.5), additional ablation studies on PrimX initialization (Sec. A.2.7), VAE designs (Sec. A.2.8), PBR extraction (Sec. A.2.9), differentiability (Sec. A.2.10) and more qualitative results (Sec. A.2.11).
  • Sec. A.3 documents the implementations details of 3DTopia-XL, including dataset and PrimX hyperparameters (Sec. A.3.1), conditioner and captions (Sec. A.3.2), model details and hyperparameters (Sec. A.3.3), and algorithms of reversible conversion between PrimX and mesh (Sec. A.3.4).
  • Besides, we also attach a demo video to demonstrate the key idea and qualitative results.
Table 5: Quantitative evaluations of Image-to-3D on the Objaverse dataset [11]. We evaluate the fidelity, diversity and distributional similarity of 3DTopia-XL for single image to 3D generation compared with existing baselines. The evaluation is performed on a subset of Objaverse consisting of 600 random samples. KIDmat denotes the KID measured in the material space by rendering materials as colored 2D images. FPD and KPD denote the FID and KID metrics measured on 3D points in the space of PointNet++ [47]. COV and MMD denote Coverage Score and Minimum Matching Distance measured in the space of Chamfer Distance.
[…]
LGM [56] Sparse-view Reconstruction 0.55 - 39.86 18.54 35.83 19.88
CRM [62] Sparse-view Reconstruction 1.93 - 37.09 14.37 31.83 25.12
ShapE [21] Native 3D Diffusion 11.95 - 20.27 7.55 59.67 14.98
LN3Diff [26] Native 3D Diffusion 0.89 - 29.36 11.27 33.50 22.64
Ours Native 3D Diffusion 1.02 0.69 15.74 3.59 50.31 14.63
Table 6: Quantitative evaluations of Image-to-3D on the GSO dataset [12]. We evaluate the fidelity, diversity and distributional similarity of 3DTopia-XL for single image to 3D generation compared with existing baselines. The evaluation is performed on a subset of GSO dataset consisting of 300 random samples. FPD and KPD denote the FID and KID metrics measured on 3D points in the space of PointNet++ [47]. COV and MMD denote Coverage Score and Minimum Matching Distance measured in the space of Chamfer Distance.
Method Paradigm KID (×$(\times 10^{-2})\downarrow$ FPD $\downarrow$ KPD (×$(\times 10^{-2})\downarrow$ COV $\uparrow$ MMD (×$(\times 10^{-3})\downarrow$
LGM [56] Sparse-view Reconstruction 0.94 34.44 11.12 33.11 23.32
CRM [62] Sparse-view Reconstruction 1.67 26.61 4.82 38.57 17.37
InstantMesh [68] Sparse-view Reconstruction 0.91 21.54 4.15 37.01 16.31
ShapE [21] Native 3D Diffusion 13.75 22.80 4.03 40.43 18.58
Ours Native 3D Diffusion 1.38 16.16 3.86 55.58 14.35
A.1 Discussion
A.1.1 Difference with Related Work
The core of our work is the proposed novel 3D representation, PrimX, that can model high-quality 3D shape, texture, and material in a unified and tensorial representation. It is worth highlighting the advantages of PrimX compared with other 3D representations in the generative context.
PrimX v.s. Implicit Vector Set.
Previous works [74, 77] introduce the implicit vector set to encode a 3D shape globally. PrimX differentiates itself from the implicit vector set in three aspects:
  • PrimX encodes not only the shape but also texture and material in a unified way, which removes the necessity for a two-stage framework [77] that generates shape and texture separately. Although the implicit vector set has the potential to model more information other than the occupancy field, there is no existing work to demonstrate and justify this design. We suspect that texture and material information are surface-aligned which is far more expensive and difficult to model for the implicit vector set that uses dense modeling of the entire 3D space by using 3D points as queries.
  • PrimX is differentiable renderable while implicit vector set can be only exported to meshes.
  • PrimX is explicit and explainable for each token feature which facilitates 1) data augmentation by applying color transformation similar to [22]; and 2) downstream tasks like inpainting by explicitly masking certain tokens.
PrimX v.s. PrimDiffusion [8].
Chen et al. [8] proposes using volumetric primitives for 3D human generation, learning primitive-based representation from multiview posed images. PrimX has unique differences:
  • PrimX requires no 3D template for generative modeling by directly denoising the position of primitives. In contrast, PrimDiffusion requires a template mesh as the anchor for all primitives which works for 3D human generation where the target subject has a shared 3D canonical space. However, this is not the case for general objects as there is no mesh template.
  • PrimX simplifies the parameter space of volumetric primitives by using only position, a single scale, and voxelized payload, while the prior work models per-axis rotation and scale factors additionally. This simplification significantly saves the computational cost while achieving a comparable quality in our preliminary study.
  • PrimX models the target’s geometry as SDF field and is capable of learning from both 3D data and 2D data. The work above models the target’s geometry as volumetric opacity field and can only learn from 2D images.
PrimX v.s. M-SDF [72].
M-SDF introduces a shape-only representation to encode SDF of 3D mesh into mosaic voxels. PrimX has two distinct differences compared to it:
  • M-SDF only represents 3D shape, while our method finds a unified way to encode shape, texture, and material with high quality. It is non-trivial to represent shape, texture, and material within our tensorial and sparse representation. The shape is typically a 3D volumetric function while textural information is only surface-aligned. It is important to note that 1) proper instantiation of texture sampling function (Sec. A.3.1) and 2) carefully designed initialization strategy (Alg. 1) for PrimX are critical for representing shape, texture and material in high quality.
  • M-SDF is specialized to 3D domain while our representation can be differentiably rendered into 2D images.
PrimX v.s. 3DGS [23].
As a trending representation for 3D reconstruction, 3DGS is known for its efficiency as a primitive-based volumetric representation. However, the number of Gaussians required to represent a high-quality 3D object is considerably high (hundreds of thousands) compared with PrimX (N=2048). This long context property will lead to training difficulty and inefficient attention computation in the generative context where the set of Gaussians is operated by DiT [45]. Instead, PrimX can be treated as an “interpolation” between fully point-based representation (3DGS) and fully voxel-based representation (dense voxel) that groups primitives into explicitly structured local voxels. This hybrid operation significantly reduces the number of primitives, leading to a shorter context that boosts the training of the Transformer.
Figure 8: 3DTopia-XL can generate 3D assets directly from single-view images without relying on 2D image-to-multiview diffusion models. We visualize the input single-view image and corresponding HDRIs for environmental lighting on the left. Please note the high-quality results and spatially varied materials generated by our method. All scenes are rendered using Blender [10].
Figure 9: 3DTopia-XL can generate 3D assets directly from texts without relying on 2D text-to-image diffusion models, which is uniquely different from sparse-view reconstruction models [19]. We visualize the input text prompts and corresponding HDRIs for environmental lighting on the left. Please note the sampling diversity and spatially varied materials generated by our method. All scenes are rendered using Blender [10].
A.1.2 Limitations and Future Work
It is important to note that 3DTopia-XL has been trained on a considerably large-scale dataset. However, there is still room for improvement in terms of quality. Different from existing high-quality 3D diffusion models [72, 77] which operate on 3D representations that are not differentiably renderable, 3DTopia-XL maintains the ability to directly learn from 2D image collections thanks to PrimX’s capability of differentiable rendering (Eq. 7 in the main paper). This opens up new opportunities to learn 3D generative models from a mixture of 3D and 2D data, which can be a solution to the lack of high-quality 3D data. Moreover, as an explicit representation, PrimX is interpretable and easy to drive. By manipulating primitives or groups of primitives, it is also fruitful to explore dynamic object generation and generative editing.
A.2 Additional Experiments
A.2.1 Quantitative Comparisons on Objaverse and GSO
We conduct extensive quantitative evaluations for image-to-3D on Objaverse [11] and GSO [12] datasets against existing methods including: 1) sparse-view reconstruction model: LGM [56], CRM [62], InstantMesh [68] and 2) native 3D diffusion models: ShapE [21] and LN3Diff [26].
Evaluation Metrics. For photometric quality, we benchmark KID [1] over the 2D renderings under random environmental lighting against ground truth. KIDmat is also calculated to measure the quality of the generated PBR materials. We render the metallic and roughness into colored 2D images according to gltf 2.0 specifics [^1] https://registry.khronos.org/glTF/specs/2.0/glTF-2.0.html#metallic-roughness-material, and measure KID against ground truth. For geometric quality, we measure Point Cloud FID [15] (FPD) and Point Cloud KID [1] using the pretrained PointNet++ [47] provided by following previous work [21, 72]. Moreover, we also evaluate Coverage Score (COV) and Minimum Matching Distance (MMD) in the space of Chamfer Distance (CD) following previous work [72, 74]. We perform the farthest point sampling over the output mesh for each method to obtain 4096 points for evaluation. We randomly sample 300 objects on the GSO dataset and 600 objects on the Objaverse dataset for evaluation, respectively.
Results. As shown in Figure 5 and Figure 6, our method achieves the best 3D geometric quality, indicating the superiority of the proposed representation and generative modeling. Note that the methods based on sparse-view reconstruction rely on pretrained 2D multiview diffusion models [52, 59] to reconstruct 3D objects from sparse input views. Therefore, their models have more input visual information and thus achieve slightly better visual quality. However, this cascaded pipeline prone to yielding 3D distortions due to 3D inconsistency of 2D diffusion models, indicated by the worse 3D metrics of them. Most importantly, 3DTopia-XL is the only method capable of producing PBR materials from images or texts among all methods.
A.2.2 Additional Comparisons
We show image-to-3D comparisons with Unique3D [64] in Fig. 10(c). As a per-sample optimization-based method using multiview images, Unique3D excels in texture quality but suffers from geometry artifacts (weird geometry from novel views) due to the inconsistency of 2D diffusion models. We believe that training a feedforward reconstruction method with PrimX would be a good geometry initialization and mesh constraint for Unique3D, which shortens its optimization time.
Moreover, we demonstrate our image-to-3D generation using a casual phone capture in Fig. 10(b), indicating the generalizability of 3DTopia-XL to the real-world domain.
Figure 10: (a) Thin structures. (b) Real-world inputs. (c) Comparison with Unique3D, which shows geometry inconsistency.
A.2.3 User Study
We conduct an extensive user study to evaluate image-to-3D performance quantitatively. We opt for an output evaluation [3] for user study, where each volunteer is shown with a pair of results comparing a random method against ours, and asked to choose the better one in four aspects: 1) Overall Quality, 2) Image Alignment, 3) Surface Smoothness, and 4) Physical Correctness. One of the samples presented to the attendees is shown in Figure 12. A total number of 48 paired samples are provided to 27 volunteers for the flip test. We summarize the average preference percentage across all four dimensions in Figure 11. 3DTopia-XL is the best one among all methods. Although the image alignment of our method is only a slight improvement against reconstruction-based methods like CRM, the superior quality of geometry and the ability to model physically based materials are the keys to producing the best overall quality in the final rendering.
Figure 11: User study. We quantitatively evaluate comparison methods by conducting preference tests against our method on four dimensions. The results show that 3DTopia-XL has the highest preference rate compared with other methods.
Figure 12: User study sample. For each sample in the user study, we present to the attendee with the input image (upper left) and target environment illuminations (bottom left) for rendering the mesh. Each volunteer is asked to choose the better one from A/B across four dimensions: 1) Overall quality, 2) Image alignment, 3) Surface smoothness, and 4) Physical correctness of renderings. The order and notation of methods are randomized and anonymized.
A.2.4 Scaling
@@
We further investigate the scaling law of 3DTopia-XL against model sizes and iterations. For metrics, we use Fréchet Inception Distance (FID) computed over 5k random samples without CFG guidance. Specifically, we consider Latent-FID which is computed in the latent space of our VAE and Rendering-FID which is computed on the DINO [44] embeddings extracted from images rendered with Eq. 7 in the main paper. Figure 13 shows how Latent-FID and Rendering-FID change as the model size increases. We observe consistent improvements as the model becomes deeper and wider. Table 7 also demonstrates that longer sequence (smaller patches) leads to better performance, which may come from the findings in the vanilla DiT that increasing GFlops leads to better performance.
Figure 13: Scaling up 3DTopia-XL improves FID. As the computation and model size scale up, the model performance improves consistently. For metrics, we consider Latent-FID which is computed in the latent space of our VAE and Rendering-FID which is computed on the DINO [44] embeddings extracted from images.
Setting Rendering-FID $\downarrow$ Latent-FID $\downarrow$
$N=256$ 76.31 104.8
$N=2048$ 16.16 24.43
A.2.5 Sampling Diversity